# 3-D CMZ I: Central Molecular Zone Overview

### v1

@@ 2 Data
Figure 1: A Herschel view of the inner 7 of the Galaxy, Red: 350 $\mu$m, Green: 160 $\mu$m, Blue: 70 $\mu$m. The main CMZ is clearly apparent by the bright concentration of emission within the inner 4 longitude and 1 latitutde, but there are additional features, such as Bania’s Clump 2 at $\ell$ $\sim$ 3.2 at higher longitudes. The twisted inner ring (Molinari et al., 2011c) is apparent in the inner 2 longitude, and some flaring up to higher latitudes is seen outside this.
@@ 3.1.1 Overview of method
Our comparatively simple approach (fitting only the cold dust component with a modified blackbody) is driven both by the availability of suitable datasets across the inner 40 of the Galaxy as well as by the fact that the intrinsic uncertainties of this analysis (such as the dust opacity, assuming only one temperature component, see more in Section 3.1.4) outweigh our measurement uncertainties. While it is possible to reduce uncertainties by attending to the detailed properties of a specific region, our approach to solve for the properties of the entire inner 40 of the Galaxy negates this possibility on a first pass. Therefore, considering the high degree of intrinsic uncertainty to any fitting approach, we aim to apply a relatively simple fitting method over a very large area, and refer the reader to more detailed and complex approaches in targeted regions (e.g. Schmiedeke et al., 2016; Marsh et al., 2015; Etxaluze et al., 2011).
At these wavelengths, the contribution of Galactic cirrus emission is significant and removal of this diffuse fore/background component is necessary. The Galactic disk abounds with low density gas and dust between the stars (Tielens, 2010; Draine, 2011; Klessen & Glover, 2016). Far-IR radiation is emitted by the dust grains in this diffuse, high-Galactic latitude medium, known as the Milky Way cirrus. This cirrus emission was first identified by IRAS (Hauser et al., 1984; Low et al., 1984), and shortly thereafter, was seen to be strongly correlated with the atomic gas column density (Boulanger & Perault, 1988) and heated by the general interstellar radiation field. This cirrus emission has been observed and analyzed in many studies since (e.g. Bianchi et al., 2017; Planck Collaboration et al., 2014). This emission is seen in Herschel wavebands (e.g. Martin et al., 2010) and in this work is excluded from the source masks and analyzed separately. For our analysis of the cirrus emission (Section 3.1.2) we include the 70 $\mu$m data point, however, for the analysis of structures within the source masks, this data point is excluded as it may be contaminated by the emission from very small dust grains (e.g. Desert et al., 1990; Compiègne et al., 2010). Additionally, as described in detail in Paper IV (Lipman et al., submitted) 70 $\mu$m is seen in extinction toward a number of regions in the CMZ and would not be suitable as a measure of the dust emission.
We perform the cirrus subtraction and modified blackbody fitting using the automated methods described in Battersby et al. (2011) with one minor change to use an automated threshold for the number of iterations (detailed below). The methods are described briefly below, but we refer the reader to Battersby et al. (2011) for a more complete description of the method, including figures. The first step is to convolve and re-grid all of the images to a common resolution (36′′), pixel-scale, and flux scale (MJy/sr).
@@ 3.1.2 Source Masks and Cirrus Emission
One of the first steps is to identify regions of significant emission associated with dense clumps and cores (we term these our ‘source masks’) rather than Galactic cirrus. The source masks and cirrus emission are identified through an iterative, automated routine that defines progressively more complete source masks with each iteration. This process is adapted from Battersby et al. (2011) and is described in full in that paper (Section 3.1, Figures 2-5). We summarize the overall procedure below, as well as a few minor updates. As in Battersby et al. (2011), we identify regions of significant and contiguous emission (“source masks”) at 500 $\mu$m where the contamination from cirrus emission decreases (Gautier et al., 1992; Peretto et al., 2010). The final cirrus emission map is the remaining emission outside the final source masks smoothed to large scales.
@@
Figure 2: The Central Molecular Zone (CMZ) contains an abundance of dense molecular gas that is highly asymmetric. This figure presents the CMZ as seen in a three-color compilation (top), in column density (middle), and dust temperature (bottom) from this work. Top: Red is N(H2) (this work), Green: 70 $\mu$(Hi-GAL; Molinari et al., 2011c, a), Blue: 8 $\mu$(GLIMPSE; Benjamin et al., 2003). Middle: Map of the column density derived in this work over the CMZ shows a high degree of asymmetry, with the majority of dense gas at positive longitudes (see §5.1) and a majority of mid-IR emission is at negative longitudes. We also note the somewhat flatter scale height of dense gas in the inner degree that fans out at higher longitudes. Bottom: Map of the dust temperature derived in this work over the CMZ show generally low dust temperatures of about 15-20 K, with a peak near the center of the Galaxy with an interesting dust ridge morphology discussed more in §5.4 and then decreasing towards positive longitudes. The temperature is generally lower at higher column densities (see Figure 6). These maps are subject to a number of uncertainties and particular care should be taken in the interpretation of points near the edges of source masks (as described in §3.1.4). Contours on the middle and bottom plots indicate N(H2) = [1,5,10] $\times$ 1022 cm-2.
@@ 3.1.3 Modified Blackbody Fits within the Source Masks
and where $\mu_{\rm H_{2}}$ is the mean molecular weight for which we adopt a value of $\mu_{\rm H_{2}}$ = 2.8 (Kauffmann et al., 2008), mH is the mass of a hydrogen atom, N(H2) is the H2 column density, $\kappa_{\nu}$ is the dust opacity, and $\gamma$ is the standard gas-to-dust ratio of 100. As in Battersby et al. (2011), we derive a dust opacity as a continuous function of frequency by fitting a power-law to the tabulated Ossenkopf & Henning (1994) dust opacities over the relevant range of frequencies. The Ossenkopf & Henning (1994) opacities assume an MRN (Mathis, Rumpl, and Nordsieck) grain size distribution (Mathis et al., 1977). Specifically, we use the version of the Ossenkopf & Henning model that assumes that the grains have thin ice mantles and have undergone 105 years of coagulation at a density of 106 cm-3, which is a reasonable and common assumption for these cold, dense clumps. This yields an expression for the dust opacity of $\kappa_{\nu}=\kappa_{0}(\nu/\nu_{0})^{\beta}$where $\kappa_{0}$ is 4.0 cm2 g-1 for $\nu_{0}=505$ GHz with a fixed $\beta$ equal to 1.75, between the standard values of 1.5-2. For this analysis we opted for a fixed $\beta$ considering the limited data available and uncertainties (see Section 3.1.4), but we direct the reader to Tang et al. (2021a) for work suggesting a dust spectral index of between $2-2.4$.
We used the package montage [^2] http://montage.ipac.caltech.edu/ to stitch and re-grid together images from the original Hi-GAL released 2×\!$\times$2 tiles. For most of the analysis, we use a mosaicked image containing the inner 7 degrees (in longitude, the latitude extent is the Hi-GAL coverage from $b=-1$ to +1) of the Galaxy, from $\ell$ = +3.5 to -3.5. This means that the automated cirrus fore/background identification and subtraction was performed over this entire region. We also mosaicked together the inner 40 degrees in longitude of the Galaxy from $\ell$ = +20 to -20, and performed cirrus fore/background identification and subtraction on this region. This complete inner 40 map is used in Figure 4, while the inner 7 map is used in the rest of the analysis. The resulting source-identified regions and column densities unsurprisingly show slight variations between the two maps in the overlapping region, due to slightly different automated fore/background identification, but are in good overall agreement.
@@ 3.1.4 Uncertainties in the Temperatures and Column Densities
The formal statistical errors on the derived dust temperatures and column densities are strong functions of the parameter values, with a median temperature error of 11% and a median column density error of 25%. These are the errors in the derived properties assuming 20% calibration uncertainty in the Hi-GAL fluxes (Molinari et al., 2010). However, the intrinsic and systematic errors are much higher. The definition and subtraction of the cirrus emission is a unique source of error for this analysis of Herschel data, since its structure and magnitude are not known a priori. Other major sources of uncertainty are the assumptions of the dust opacity and the gas to dust ratio. The gas to dust ratio may be lower if there is a higher metallicity at the Galactic Center, which would affect our mass and density estimates linearly. Based on the Monte-Carlo error simulations of Battersby et al. (2010) using similar analyses, we expect that our measured column densities have a systematic uncertainty of about a factor of two. The relative uncertainty between regions in our maps, however, is smaller, since many of these systematic uncertainties (such as the choice of gas to dust ratio) apply equally to all regions.
@@ 3.1.5 Data Release
We release all the analysis products from this work (including fore/background-subtracted column density and dust temperature maps over the inner 40 of the Galaxy) in the 3-D CMZ Harvard Dataverse: https://dataverse.harvard.edu/dataverse/3D_CMZ. The Harvard Dataverse is an online data repository that enables long-term data preservation and sharing. The details of the data products available are described in Appendix A as well as in the online meta-data. Project updates can also be found on the 3-D CMZ website: https://centralmolecularzone.github.io/3D_CMZ/.
@@ 4.1 Column Density in the CMZ
The column density map over the inner 3.5 longitude of the Galaxy is shown in Figure 2. The peak column density in the map is toward the SgrB2 complex with a value of about 2 $\times$ 1024 cm-2. However, the peak density is observed towards several pixels which are saturated in the Herschel data towards Sgr B2 North and Main, indicating a saturated pixels towards very bright emission. Missing this bright emission (likely associated with very high column density gas) and beam-smearing of the relatively coarse Herschel data both act to lower our measurement and the true column density in this region is certainly higher (as demonstrated with higher resolution data, e.g. Hatchfield et al., 2020; Ginsburg et al., 2018; Schmiedeke et al., 2016). Typically “clouds” in the CMZ have edges at about N(H2) $\sim$1023 cm-2 (this threshold was used to define the boundary of CMZ clouds in the CMZoom survey; see Battersby et al., 2020) with peaks of a few times 1023 cm-2 in their centers.
These column densities are sufficient to absorb a substantial fraction of the soft (E $<$ 4.5 keV) X-ray emission in the Galactic center, as can be seen in deep XMM-Newton images Ponti et al. (2015), where clouds like Sgr B2, the Brick, and Sgr C appear in silhouette. The CMZ is unique in our galaxy for having an extremely large molecular gas fraction ($N_{H2}/N_{H}>90$%; Sofue, 2022). As a result, the Compton thickness of clouds in this region is most correctly calculated relative to the column density of molecular hydrogen. Taking the cross section of H2 to be 2.8 larger than H at high energies (Yan et al., 1998), CMZ clouds would become Compton thick for $N$(H2) $>$5.4 ×$\times 10^{23}$ cm-2. While only the Sgr B2 cloud exceeds this limit in our Galaxy’s center at these resolutions, it is likely higher on smaller spatial scales, and a number of other regions are within a factor of two (our estimated uncertainty) of this value. This demonstrates the importance that gas in the central $R\sim 200$ of galaxies could more generally play in obscuring extragalactic AGN.
@@ 4.2 Dust Temperature in the CMZ
The dust temperature map in the bottom panel of Figure 2 shows the range in dust temperatures derived in this work [^3] There exist spurious pixels near the source edges with temperatures ranging from 5 to 45 K, but the majority are within the stated range, from about 12 K to about 35 K with a median dust temperature of 21 K and a standard deviation of 4 K. The highest dust temperatures are in the very central regions of the CMZ, but avoid the dense cloud regions.
@@ 4.3 The Inner 40 of the Galaxy
We note here that while the fore/background subtraction algorithm was the same for the inner 7 maps and the inner 40 maps, the application across the different mosaicked fields results in slightly different source identification, and therefore also different fore/background levels and values of the column density and dust temperature in the overlap region. These differences are well within the quoted uncertainties (see Section 3.1.4). We use the inner 7 maps for all of the remaining analysis in this work.
@@ 5.2 Insights on the 3-D CMZ from Dust Continuum: Extent
Inner 7$\times$ 2555All ‘inner’ values are centered on [$\ell$,b]=[0,0] 1000 $\times$ 280 2.8 ×$\times 10^{7}$ (7.8 ×$\times 10^{7}$) 74% (58%) 26% (42%)
@@
Inner 1.05$\times$ 1.05666To match the region from Ferrière et al. (2007), who found a total mass of about 1.3 ×$\times 10^{7}$ M 150 $\times$ 150 0.6 ×$\times 10^{7}$ (1.0 ×$\times 10^{7}$) 56% (52%) 44% (48%)
$\ell$ = -1.05 to 2.25, b = $\pm$ 0.75777To match the region from Dahmen et al. (1998), who found a total mass of $\sim 1.2-6.4\times 10^{7}$ M 470 $\times$ 210 2.3 ×$\times 10^{7}$ (4.1 ×$\times 10^{7}$)
$\ell$ = -1.0 to 1.7, b = $\pm$ 0.5888To match the Henshaw et al. (2023) definition of the CMZ region 380 $\times$ 140 2.0 ×$\times 10^{7}$ (3.1 ×$\times 10^{7}$)
$\ell$ = -1.3 to 1.8, b = $\pm$ 0.75999Extent measured from this work 440 $\times$ 210 2.2 ×$\times 10^{7}$ (3.9 ×$\times 10^{7}$)

### v2

@@ 2 Data
Figure 1: A Herschel view of the inner 7 of the Galaxy, Red: 350 $\mu$m, Green: 160 $\mu$m, Blue: 70 $\mu$m. The main CMZ is clearly apparent by the bright concentration of emission within the inner 4 longitude and 1 latitude, but there are additional features, such as Bania’s Clump 2 at $\ell$ $\sim$ 3.2 at higher longitudes. The twisted inner ring (Molinari et al., 2011c) is apparent in the inner 2 longitude, and some flaring up to higher latitudes is seen outside this.
@@ 3.1.1 Overview of method
Our comparatively simple approach is driven both by the availability of suitable datasets across the inner 40 of the Galaxy as well as by the fact that the intrinsic uncertainties of this analysis (such as the dust opacity, assuming only one temperature component, see more in Section 3.1.4) outweigh our measurement uncertainties. While it is possible to reduce uncertainties by attending to the detailed properties of a specific region, our approach to solve for the properties of the entire inner 40 of the Galaxy negates this possibility on a first pass. Therefore, considering the high degree of intrinsic uncertainty to any fitting approach, we aim to apply a relatively simple fitting method over a very large area, and refer the reader to more detailed and complex approaches in targeted regions (e.g. Schmiedeke et al., 2016; Marsh et al., 2015; Etxaluze et al., 2011).
At these wavelengths, the contribution of Galactic cirrus emission is significant and removal of this diffuse fore/background component is necessary. We perform the cirrus subtraction and modified blackbody fitting using the automated methods described in Battersby et al. (2011) with one minor change to use an automated threshold for the number of iterations (detailed below). The methods are described briefly below, but we refer the reader to Battersby et al. (2011) for a more complete description of the method, including figures. The first step is to convolve and re-grid all of the images to a common resolution (36′′), pixel-scale, and flux scale (MJy/sr).
@@ 3.1.2 Source Masks and Cirrus Emission
One of the first steps is to identify regions of significant emission associated with dense clumps and cores (we term these our ‘source masks’) rather than Galactic cirrus. The Galactic disk abounds with low density gas and dust between the stars (Tielens, 2010; Draine, 2011; Klessen & Glover, 2016). Far-IR radiation is emitted by the dust grains in this diffuse, high-Galactic latitude medium, known as the Milky Way cirrus. This cirrus emission was first identified by IRAS (Hauser et al., 1984; Low et al., 1984), and shortly thereafter, was seen to be strongly correlated with the atomic gas column density (Boulanger & Perault, 1988) and heated by the general interstellar radiation field. This cirrus emission has been observed and analyzed in many studies since (e.g. Bianchi et al., 2017; Planck Collaboration et al., 2014). This emission is seen in Herschel wavebands (e.g. Martin et al., 2010) and in this work is excluded from the source masks and analyzed separately. For our analysis of the cirrus emission (Section 3.1.2) we include the 70 $\mu$m data point, however, for the analysis of structures within the source masks, this data point is excluded as it may be contaminated by the emission from very small dust grains (e.g. Desert et al., 1990; Compiègne et al., 2010). Additionally, as described in detail in Paper IV (Lipman et al., submitted) 70 $\mu$m is seen in extinction toward a number of regions in the CMZ and would not be suitable as a measure of the dust emission.
The source masks and cirrus emission are identified through an iterative, automated routine that defines progressively more complete source masks with each iteration. This process is adapted from Battersby et al. (2011) and is described in full in that paper (Section 3.1, Figures 2-5). We summarize the overall procedure below, as well as a few minor updates. As in Battersby et al. (2011), we identify regions of significant and contiguous emission (“source masks”) at 500 $\mu$m where the contamination from cirrus emission decreases (Gautier et al., 1992; Peretto et al., 2010). The final cirrus emission map is the remaining emission outside the final source masks smoothed to large scales.
@@
Figure 2: The Central Molecular Zone (CMZ) contains an abundance of dense molecular gas that is highly asymmetric. This figure presents the CMZ as seen in a three-color compilation (top), in column density (middle), and dust temperature (bottom) from this work. Top: Red is N(H2) (this work), Green: 70 $\mu$(Hi-GAL; Molinari et al., 2011c, a), Blue: 8 $\mu$(GLIMPSE; Benjamin et al., 2003). Middle: Map of the column density derived in this work over the CMZ shows a high degree of asymmetry, with the majority of dense gas at positive longitudes (see §5.1) and a majority of mid-IR emission is at negative longitudes. We also note the somewhat flatter scale height of dense gas in the inner degree that fans out at higher longitudes. Bottom: Map of the dust temperature derived in this work over the CMZ show generally low dust temperatures of about 15-20 K, with a peak near the center of the Galaxy with an interesting dust ridge morphology discussed more in §5.4 and then decreasing towards positive longitudes. The temperature is generally lower at higher column densities (see Figure 6). These maps are subject to a number of uncertainties and particular care should be taken in the interpretation of points near the edges of source masks (as described in §3.1.4). Contours on the middle and bottom plots indicate N(H2) = [1,5,10] $\times$ 1022 cm-2. The resolution of the column density and dust temperature maps are 36′′ (or 1.4 pc at the assumed CMZ distance) as described in Section 3.1.1
@@ 3.1.3 Modified Blackbody Fits within the Source Masks
The variables in equation 1 are: $S_{\nu}$ is the flux density as a function of frequency, $\nu$, in units of Watts per steradian per square meters per Hz. $h$ is Planck’s constant, $c$ is the speed of light, $k$ is the Boltzmann constant, and $T$ is the temperature of the material. $\tau_{\nu}$ is defined by Equation 2 where $\mu_{\rm H_{2}}$ is the mean molecular weight for which we adopt a value of $\mu_{\rm H_{2}}$ = 2.8 (Kauffmann et al., 2008), mH is the mass of a hydrogen atom, N(H2) is the H2 column density, $\kappa_{\nu}$ is the dust opacity, and $\gamma$ is the gas-to-dust ratio. For the gas to dust ratio, we adopt a standard value of 100. While this gas to dust ratio is commonly used (e.g. Barnes et al., 2017; Battersby et al., 2010, 2011), it is a largely unknown quantity, without even clear uncertainties on the value (e.g. Lv et al., 2017), see Section 3.1.4 for more details). As in Battersby et al. (2011), we derive a dust opacity as a continuous function of frequency by fitting a power-law to the tabulated Ossenkopf & Henning (1994) dust opacities over the relevant range of frequencies. The Ossenkopf & Henning (1994) opacities assume an MRN (Mathis, Rumpl, and Nordsieck) grain size distribution (Mathis et al., 1977). Specifically, we use the version of the Ossenkopf & Henning model that assumes that the grains have thin ice mantles and have undergone 105 years of coagulation at a density of 106 cm-3, which is a reasonable and common assumption for these cold, dense clumps. This yields an expression for the dust opacity of $\kappa_{\nu}=\kappa_{0}(\nu/\nu_{0})^{\beta}$where $\kappa_{0}$ is 4.0 cm2 g-1 for $\nu_{0}=505$ GHz with a fixed $\beta$ equal to 1.75, between the standard values of 1.5-2. For this analysis we opted for a fixed $\beta$ considering the limited data available and uncertainties (see Section 3.1.4), but we direct the reader to Tang et al. (2021a) for work suggesting a dust spectral index of between $2-2.4$.
We used the package montage [^2] http://montage.ipac.caltech.edu/ to stitch and re-grid together images from the original Hi-GAL released 2×\!$\times$2 tiles. For most of the analysis, we use a mosaicked image containing the inner 7 degrees (in longitude, the latitude extent is the Hi-GAL coverage from $b=-1$ to +1) of the Galaxy, from $\ell$ = +3.5 to -3.5. This means that the automated cirrus fore/background identification and subtraction was performed over this entire region. We also mosaicked together the inner 40 degrees in longitude of the Galaxy from $\ell$ = +20 to -20, and performed cirrus fore/background identification and subtraction on this region. This complete inner 40 map is used in Figure 4, while the inner 7 map is used in the rest of the analysis. The resulting source-identified regions and column densities unsurprisingly show slight variations between the two maps in the overlapping region, due to slightly different automated fore/background identification, however, these differences are well within the quoted uncertainties (see Section 3.1.4).
@@ 3.1.4 Uncertainties in the Temperatures and Column Densities
The formal statistical errors on the derived dust temperatures and column densities are strong functions of the parameter values, with a median temperature error of 11% and a median column density error of 25%. These are the errors in the derived properties assuming 20% calibration uncertainty in the Hi-GAL fluxes (Molinari et al., 2010). However, the intrinsic and systematic errors are much higher. The definition and subtraction of the cirrus emission is a unique source of error for this analysis of Herschel data, since its structure and magnitude are not known a priori. Other major sources of uncertainty are the assumptions of the dust opacity and the gas to dust ratio. We have very limited data on the validity of the commonly used gas to dust ratio of 100 (e.g. Lv et al., 2017). Due to the different temperatures, densities, and other physical properties in the Galactic Center (e.g. Henshaw et al., 2023), it is quite plausible that the gas to dust ratio may be different, which would have a large impact on the inferred quantities. The gas to dust ratio may be lower if there is a higher metallicity at the Galactic Center, which would affect our mass and density estimates linearly. Based on the Monte-Carlo error simulations of Battersby et al. (2010) using similar analyses, we expect that our measured column densities have a systematic uncertainty of about a factor of two. However, the Battersby et al. (2010) Monte Carlo simulations did not include all possible sources of errors, including the gas to dust ratio. So, the true systematic uncertainty may be even higher. The relative uncertainty between regions in our maps, however, is smaller, since many of these systematic uncertainties (such as the choice of gas to dust ratio) apply equally to all regions.
@@ 3.1.5 Data Release
We release all the analysis products from this work (including fore/background-subtracted column density and dust temperature maps over the inner 40 of the Galaxy) in the 3-D CMZ Harvard Dataverse: https://dataverse.harvard.edu/dataverse/3D_CMZ. The dataset for this paper is found in the Papers I and II repository: https://doi.org/10.7910/DVN/7DOJG5 (Battersby et al., 2024). The Harvard Dataverse is an online data repository that enables long-term data preservation and sharing. The details of the data products available are described in Appendix A as well as in the online meta-data. Project updates can also be found on the 3-D CMZ website: https://centralmolecularzone.github.io/3D_CMZ/.
@@ 4.1 Column Density in the CMZ
The column density map over the inner 3.5 longitude of the Galaxy is shown in Figure 2. The peak column density in the map is toward the SgrB2 complex with a value of about 2 $\times$ 1024 cm-2. However, the peak density is observed towards several pixels which are saturated in the Herschel data towards Sgr B2 North and Main  [^3] Sgr B2 North and Main refer to the densest portions of the Sgr B2 molecular cloud, marked in Figure 1, and have coordinates of approximately $\alpha_{J2000}=17h47m19.88s$, $\delta_{J2000}=-28^{\circ}22^{\prime}18.4^{\prime\prime}$ and $\alpha_{J2000}=17h47m20.35s$, $\delta_{J2000}=-28^{\circ}23^{\prime}3.0^{\prime\prime}$, respectively, please see Schmiedeke et al. (2016) for more details., indicating a saturated pixels towards very bright emission. Missing this bright emission (likely associated with very high column density gas) and beam-smearing of the relatively coarse Herschel data both act to lower our measurement and the true column density in this region is certainly higher (as demonstrated with higher resolution data, e.g. Hatchfield et al., 2020; Ginsburg et al., 2018; Schmiedeke et al., 2016). Typically “clouds” in the CMZ have edges at about N(H2) $\sim$1023 cm-2 (this threshold was used to define the boundary of CMZ clouds in the CMZoom survey; see Battersby et al., 2020) with peaks of a few times 1023 cm-2 in their centers.
These column densities are sufficient to absorb a substantial fraction of the soft (E $<$ 4.5 keV) X-ray emission in the Galactic center, as can be seen in deep XMM-Newton images Ponti et al. (2015), where clouds like Sgr B2, the Brick, and Sgr C appear in silhouette. The CMZ is unique in our galaxy for having an extremely large molecular gas fraction ($N_{H2}/N_{H}>90$%; Sofue, 2022). As a result, the Compton thickness of clouds in this region is most correctly calculated relative to the column density of molecular hydrogen. Taking the cross section of H2 to be 2.8 larger than H at high energies (Yan et al., 1998), CMZ clouds would become Compton thick for $N$(H2) $>$5.4 ×$\times 10^{23}$ cm-2. While only the Sgr B2 cloud exceeds this limit in our Galaxy’s center at these resolutions, it is likely higher on smaller spatial scales, and a number of other regions are within a factor of two (our estimated uncertainty) of this value. Other galaxies may have even denser central regions (e.g. Stuber et al., 2023). If a substantial fraction of the gas exceeds these limits, and if the viewing angle towards a central supermassive black hole is obscured by this dense gas, then gas in the central $R\sim 200$ pc of galaxies could play in obscuring extragalactic AGN.
@@ 4.2 Dust Temperature in the CMZ
The dust temperature map in the bottom panel of Figure 2 shows the range in dust temperatures derived in this work [^4] There exist spurious pixels near the source edges with temperatures ranging from 5 to 45 K, but the majority (99.5%) are within the stated range of 12 to 35 K., from about 12 K to about 35 K with a median dust temperature of 21 K and a standard deviation of 4 K. The highest dust temperatures are in the very central regions of the CMZ, but avoid the dense cloud regions.
@@ 5.2 Insights on the 3-D CMZ from Dust Continuum: Extent
Inner 7$\times$ 2666All ‘inner’ values are centered on [$\ell$,b]=[0,0] 1000 $\times$ 280 2.8 ×$\times 10^{7}$ (7.8 ×$\times 10^{7}$) 74% (58%) 26% (42%)
@@
Inner 1.05$\times$ 1.05777To match the region from Ferrière et al. (2007), who found a total mass of about 1.3 ×$\times 10^{7}$ M 150 $\times$ 150 0.6 ×$\times 10^{7}$ (1.0 ×$\times 10^{7}$) 56% (52%) 44% (48%)
$\ell$ = -1.05 to 2.25, b = $\pm$ 0.75888To match the region from Dahmen et al. (1998), who found a total mass of $\sim 1.2-6.4\times 10^{7}$ M 470 $\times$ 210 2.3 ×$\times 10^{7}$ (4.1 ×$\times 10^{7}$)
$\ell$ = -1.0 to 1.7, b = $\pm$ 0.5999To match the Henshaw et al. (2023) definition of the CMZ region 380 $\times$ 140 2.0 ×$\times 10^{7}$ (3.1 ×$\times 10^{7}$)
$\ell$ = -1.3 to 1.8, b = $\pm$ 0.75101010Extent measured from this work 440 $\times$ 210 2.2 ×$\times 10^{7}$ (3.9 ×$\times 10^{7}$)